the present age of a father is three years more than three times the age of his son. three year hence,fathers age will be 10 years more than twice the age of his son. determine their present ages.
Answer by ankor@dixie-net.com(22740)

Let f = father's age; let s = son's age
:
"the present age of a father is three years more than three times the age of his son."
f = 3s + 3
:
"three year hence, fathers age will be 10 years more than twice the age of his son."
f + 3 = 2(s+3) + 10
f + 3 = 2s + 6 + 10
f + 3 = 2s + 16
f = 2s + 16 - 3
f = 2s + 13
:
determine their present ages.
Substitute (3s+3) for f in the above equation, find s
3s+3 = 2s + 13
3s - 2s = 13 - 3
s = 10 yrs, son's present age
then
f = 3(10) + 3
f = 33 yrs, father's present age
;
:
Check solution in the statement:
"three year hence,fathers age will be 10 years more than twice the age of his son.
33 + 3 = 2(10+3) + 10
36 = 2(13) + 10